# weaker gravitational waves

### v1

@@ Abstract
We study the evolution of domain wall networks and their phenomenological implications in a model of a real scalar $\chi$, where a $Z_{2}$-symmetry is slightly broken by a potential bias $V_{bias}$. It is demonstrated that the latter triggers domain wall annihilation considerably earlier than previously thought. Namely, we observe that the scaling relation $t_{ann}\propto 1/V^{2/3}_{bias}$ for the annihilation time $t_{ann}$ fits to the simulation data better than a commonly assumed $t_{ann}\propto 1/V_{bias}$. As a result, the energy density of gravitational waves produced by the network of biased domain walls, for a given tiny $V_{bias}$, is suppressed compared to naive expectations. The spectral shape of gravitational waves is similar to that resulting from unbiased domain walls, but with more power in the close-to-maximum ultraviolet part. In the far ultraviolet region, the spectrum of gravitational waves becomes nearly flat; such a plateau has been recognized earlier in the case of unbiased walls. In our investigation we mainly focus on the symmetry breaking potential $V_{breaking}\propto\chi^{3}$, and argue that no significant modifications of the domain walls evolution take place if one includes higher powers of $\chi$.
@@ 1 Introduction
It is often considered that the slight symmetry breaking is irrelevant for most of DW evolution, and the network collapse happens almost instantly at a certain cosmological time $t_{ann}$. The latter is commonly inferred from the equality between the energy density of DWs assumed to be in the scaling regime and the constant potential bias $V_{bias}$ [^1] The case of time-dependent potential bias has been considered in Refs. [16, 17], and we do not discuss it here., i.e., the difference of potential energy densities of two vacuums. This leads to the following dependence of the annihilation time on the potential bias: $t_{ann}\propto 1/V_{bias}$ [8], see Sec. 2 for more details. Furthermore, one usually supposes that the potential bias has a little impact on the shape of GW power spectrum: before the annihilation, at times $t<t_{ann}$ GWs are produced as if $Z_{2}$-symmetry were exact, and at later times, $t>t_{ann}$, the GW production entirely terminated and GW spectrum evolves only due to cosmic expansion. This somewhat simplified picture came under scrutiny with high resolution simulations becoming available [16, 17, 18, 19, 20] [^2] In the early work on the subject [21], the issue with limited resolution was circumvented by assuming the wall width growing with time. This technique developed in Ref. [22] allowed for long simulation times. We do not use it in the present work..
@@
The shapes of GWs power spectrum obtained with and without potential bias [25, 26, 27] are similar, but there are notable differences. In both cases, there is a peak at frequency $f_{p}$ defined by the correlation length of the DW network, which is of the order of the Hubble radius at annihilation. The far infrared (IR) part of the spectrum is mostly determined by causality considerations rather than particularities of DW evolution with the characteristic slope approaching $\Omega_{gw}\propto f^{3}$ [28, 29] at $f\ll f_{peak}$. In the close-to-maximum ultraviolet (UV) part of the spectrum, $f\gtrsim f_{peak}$, we observe a significantly softer decrease of the spectrum compared to the unbiased case. A similar finding has been made recently in Ref. [17]. Interestingly, the plateau in the far UV region, $f\gg f_{peak}$, already present in the unbiased case [27] also occurs for biased DWs. The spectrum ends with the characteristic exponential falloff at frequencies exceeding the inverse DW width, which is also a universal feature.
@@ 2 Theoretical setup
Let us first briefly review basics of DW evolution in the limit $V_{bias}\rightarrow 0$, where the exact $Z_{2}$-symmetry is restored. We assume that the field $\chi$ is initially set to zero cosmologically and starts rolling to its potential minima after the Hubble parameter $H=\dot{a}/a$ drops enough to fulfill the condition $H\lesssim\sqrt{\lambda}v$, from which point on the DW network comes into existence. Note that we neglect any non-gravitational interactions of $\chi$ with the primordial plasma, and hence its thermal mass is set to zero. Soon after its formation, the DW network enters the scaling regime [22]. Numerical simulations [27] suggest that the scaling starts universally, i.e., independently of initial conditions, at the time $t_{sc}$, when the DW width $\delta_{wall}$, estimated as
@@
becomes small compared to the Hubble radius, i.e., $\delta_{wall}\simeq 0.05H^{-1}_{sc}$. In the scaling regime, the long-range dynamics of the DW network is defined by only one parameter, the expansion rate of the Universe, $H$. In particular, simulations reveal that at any time there is essentially one long DW with a characteristic curvature radius $H^{-1}$ in the Hubble patch. Closed DWs are also present, but their impact on dynamics is negligible [27]. As a result, the energy density of the network is estimated as
@@
From now on we switch to the case of biased DWs, unless otherwise specified. Conventionally, one defines the DW network annihilation time $\tilde{\tau}_{ann}$ (the reason for the tilde notation will become clear shortly) from the balance between the DW energy density and the potential bias, i.e., $\rho_{wall}\sim V_{bias}$. One also assumes that the scaling of DWs continues up to the annihilation time, so that $\rho_{wall}(\tilde{\tau}_{ann})\sim\sigma_{wall}H_{ann}$. This leads to the dependence $\tilde{\tau}_{ann}\propto 1/\sqrt{\epsilon}$ of so defined annihilation time $\tilde{\tau}_{ann}$ [8]. However, we observe in Sec. 4 that departures from the scaling regime start significantly earlier, thus compromising the physical interpretation of $\tilde{\tau}_{ann}$. Therefore, we consider an alternative way of defining the DW network collapse time, which will prove to be more relevant. Following Ref. [19], we introduce the false vacuum fraction:
@@
Naively one would expect that $\tau_{ann}\sim\tilde{\tau}_{ann}\propto 1/\sqrt{\epsilon}$. However, we in Sec. 4 that the simulation data favor a different behavior. Hereafter we adopt the definition of the DW annihilation time (10).
@@ 3 Numerical setup
$\chi^{\prime\prime}+\frac{2a^{\prime}}{a}\chi^{\prime}-\partial^{2}_{i}\chi+\tau^{2}\chi(\chi^{2}-1)+3\epsilon\tau^{2}\chi^{2}=0\;.$ (15)
@@
Optimally by the end of simulations, the DW width should not be smaller than the lattice spacing, i.e.,
@@
where $N=1024$ or $2048$ is the lattice grid number, $\kappa={\cal O}(1)$ is a constant, $L_{i}/N$, where $L_{i}=a(\tau_{i})L=L$ is the initial lattice spacing and we take into account that the lattice spacing grows linearly with the scale factor. We also require that the final Hubble horizon is still smaller than the simulation box, i.e.,
@@
So below we fix these final times, but allow for variations of $\kappa^{\prime}$ and $\kappa$. The major reason is that DW phenomenology and GW signal in particular are dominated by the Hubble scale physics at the latest stages of the network existence. Therefore, we wish the simulation box size to be considerably larger than the Hubble rate at those times, i.e., we allow the parameter $\kappa^{\prime}$ to be larger than unity. With such a choice, however, one runs the risk that DWs can become too thin relative to the lattice spacing by the end of simulations. As it has been observed in Ref. [27], this does not invalidate simulations, but leads to the appearance of the artificial features in the UV part of the spectrum propagating towards IR with time. The level of “contamination” can be controlled by choosing different $\kappa^{\prime}$ and $\kappa$
@@
Hereafter we mainly take $\kappa^{\prime}=\pi\kappa/2$. However, in order to reconstruct the GW spectrum, one will need to consider all three options. In this way, we can “zoom in” the IR, middle, and UV parts of the GW spectrum, and then glue all three parts. The validity of this techniques was proven in Ref. [27], where unbiased DWs have been studied by comparing results obtained with lattices of different resolution.
@@ 4 Numerical results: evolution of domain wall network
Figure 1: Snapshots of DW evolution obtained in the case of dimensionless bias parameter $\epsilon=0.01$. DWs are shown with green, while the regions filled with true and false vacuum are shown with white and red, respectively. Simulations have been carried out with $1024^{3}$ lattice.
@@
Evolution of the false vacuum fraction demonstrated in Fig. 4 is particularly illuminating. We parameterise the false vacuum fraction as in Ref. [19]:
@@
where $p$ is assumed to be a constant. See also Refs. [30, 31, 32] for theoretical approaches to the problem of DW collapse. Note that the parameter $\tau_{ann}$ entering Eq. (25) matches the definition of the DW annihilation time given in Eq. (10). Compared to Ref. [19], we observe that the power $p$ may vary as a function of the potential bias, and at this point we observe dependence on the initial conditions illustrated in Fig. 5. Namely, the variation is rather strong for $k_{cut}=1$ and $k_{cut}=5$, while for $k_{cut}=0.3$ the parameter $p$ gets confined to a narrow range near $p=3.3$. In this “stabilised” regime, values of the parameter $p$ are in a good agreement with those of Ref. [19].
@@
where $C(k_{cut})$ is a coefficient, which is independent of $\epsilon$, but slightly varying with the momentum cutoff $k_{cut}$. In particular, one has $C(1)\approx 1$, $C(0.3)\approx 1.4$, and $C(5)\approx 0.9$. Recalling that the initial time $\tau_{i}$ is chosen in such a way that $H_{i}=1/(a_{i}\tau_{i})=\sqrt{\lambda}v$, one can rephrase (26) in terms of the Hubble rate at radiation domination as
@@
$\frac{H_{ann}}{\sqrt{\lambda}v}=\frac{0.1}{C^{2}(k_{cut})}\cdot\left(\frac{\epsilon}{\lambda v}\right)^{2/3}\;.$ (27)
@@
Interestingly, the dependence $\tau_{ann}(\epsilon)$ as in Eq. (26) is consistent with the standard estimate $\rho_{wall}\sim V_{bias}$ provided that the DW energy density drops as $\rho_{wall}\propto 1/a^{3}$ rather than $\rho_{wall}\propto 1/a^{2}$. In other words, in order to satisfy the standard estimate $\rho_{wall}\sim V_{bias}$, the DW density must decrease faster than in the scenario with unbiased DWs. This correlates with the behavior of the scaling parameter shown in Fig. 3 suggesting that the potential bias starts impacting DWs early in their evolution. Such a similarity of the behavior of biased DWs with the case of non-interacting dust particles might be not a coincidence (e.g. it may be partly due to growing amount of closed DW behaving like dust inside a Hubble patch), and it is worth investigating in the future. Notably, the observed dependence $\tau_{ann}(\epsilon)$ is largely independent of initial conditions. Namely, unlike the parameter $p$ and the coefficient $C(k_{cut})$, the relation $\tau_{ann}\propto 1/\epsilon^{1/3}$ is robust against variations of the momentum cutoff $k_{cut}$. In other words, we find that DW annihilation time is parametrically smaller (with respect to $\epsilon$) than naively expected. As we see below, this leads to drastic decrease of expected GW power as compared to the commonly assumed calculations.
@@
To derive the relation (26) we have probed the ranges $\epsilon=[0.0035,0.05]$ and $\epsilon=[0.01,0.05]$ using $2048^{3}$ and $1024^{3}$ lattices, respectively. Note that the similar conclusion can be derived from the analysis of false vacuum fraction performed in Ref. [19], where the narrow range $\epsilon=[4\cdot 10^{-4},8\cdot 10^{-4}]$ has been explored with higher resolution simulations. One can check that the behavior (26) is indeed obeyed except for the smallest $\epsilon$ there. Steepening of the behaviour $\tau_{ann}(\epsilon)$ in the range of very small $\epsilon$ is likely attributed to the fact that in this case the network decay takes place at sufficiently late times, when effects, related to finite lattice size/resolution, become significant.
@@
Figure 6: Comparison of the potential bias impact on DW evolution caused by the symmetry breaking potentials $V_{breaking}=\epsilon\chi^{3}$, $\epsilon\chi v^{3}$, and $\epsilon(\chi^{5}/v^{2}+\chi^{6}/v^{3})$. The bias parameter $\epsilon$ has been set to $\epsilon=0.05$ (in units of $\lambda v$). The impact is demonstrated in terms of the scaling parameter $\xi$ (left panel) and the false vacuum fraction ${\cal F}_{fv}$ (right panel). Simulations have been carried out with $1024^{3}$ lattice.
@@
Let us consider other choices of $Z_{2}$-symmetry breaking caused by the linear term $V_{breaking}=\epsilon\chi v^{2}$ as well as the non-renormalisable term of the form $V_{breaking}=\epsilon\chi^{5}/v^{2}+\epsilon^{\prime}\chi^{6}/v^{3}$. In the latter case, the term $\epsilon^{\prime}\chi^{6}/v^{3}$ with some positive constant $\epsilon^{\prime}$ has been added to keep the overall potential bounded from below. As it is clear from Fig. 6, the impact of the higher order potential bias on DW evolution is almost indistinguishable from that of the cubic bias. On the other hand, the effect of the linear bias is dramatically different. The likely reason is that the linear bias term changes behaviour of its potential $V(\chi)$ around the minimum $\chi=0$ leading to non-zero $\partial V/\partial\chi$ at $\chi=0$. As a result, one of the minima becomes slightly preferred meaning that the linear potential bias triggers the population bias. This may explain, why the DW network is destroyed at much faster rate for the same $\epsilon$, see Ref. [33], and consequently the expected GW signal is suppressed relative to the case of the cubic bias. In this work we focus on the effects, which are entirely due to the potential bias, and omit any further discussion of the case $V_{breaking}\propto\chi$.

### v2

@@ Abstract
We study the evolution of domain wall networks and their phenomenological implications in a model of a real scalar $\chi$, where a $Z_{2}$-symmetry is slightly broken by a potential bias $V_{bias}$. It is demonstrated that the latter triggers domain wall annihilation considerably earlier than previously thought. Namely, we observe that the scaling relation $t_{ann}\propto 1/V^{2/3}_{bias}$ for the annihilation time $t_{ann}$ fits to the simulation data better than a commonly assumed $t_{ann}\propto 1/V_{bias}$. As a result, the energy density of gravitational waves produced by the network of biased domain walls, for a given tiny $V_{bias}$, is suppressed compared to naive expectations. The spectral shape of gravitational waves is similar to that resulting from unbiased domain walls, but with more power in the close-to-maximum ultraviolet part. In the far ultraviolet region, the spectrum of gravitational waves becomes nearly flat; such a plateau has been recognised earlier in the case of unbiased walls. In our investigation we mainly focus on the symmetry breaking potential $V_{breaking}\propto\chi^{3}$, and argue that no significant modifications of the domain walls evolution take place if one includes higher powers of $\chi$.
@@ 1 Introduction
It is often considered that the slight symmetry breaking is irrelevant for most of DW evolution, and the network collapse happens almost instantly at a certain cosmological time $t_{ann}$. The latter is commonly inferred from the equality between the energy density of DWs assumed to be in the scaling regime and the constant potential bias $V_{bias}$ [^1] The case of time-dependent potential bias has been considered in Refs. [16, 17, 18], and we do not discuss it here., i.e., the difference of potential energy densities of two vacuums. This leads to the following dependence of the annihilation time on the potential bias: $t_{ann}\propto 1/V_{bias}$ [8], see Sec. 2 for more details. Furthermore, one usually supposes that the potential bias has a little impact on the shape of GW power spectrum: before the annihilation, at times $t<t_{ann}$, GWs are produced as if $Z_{2}$-symmetry were exact, and at later times, $t>t_{ann}$, the GW production is entirely terminated and GWs evolve only due to cosmic expansion. This somewhat simplified picture came under scrutiny with high resolution simulations becoming available [16, 17, 18, 19, 20, 21] [^2] In the early work on the subject [22], the issue with limited resolution was circumvented by assuming the wall width growing with time. This technique developed in Ref. [23] allowed for long simulation times. We do not use it in the present work..
@@
The shapes of GW power spectra obtained with and without potential bias [26, 27, 28] are similar, but there are notable differences. In both cases, there is a peak at frequency $f_{peak}$ defined by the correlation length of the DW network, which is of the order of the Hubble radius at annihilation. The far infrared (IR) part of the spectrum is mostly determined by causality considerations rather than particularities of DW evolution with the characteristic slope approaching $\Omega_{gw}\propto f^{3}$ [29, 30] at $f\ll f_{peak}$. In the close-to-maximum ultraviolet (UV) part of the spectrum, $f\gtrsim f_{peak}$, we observe a significantly softer decrease of the spectrum compared to the unbiased case. A similar finding has been made recently in Ref. [17]. Interestingly, the plateau in the far UV region, $f\gg f_{peak}$, already present in the unbiased case [28], also occurs for biased DWs. The spectrum ends with the characteristic exponential falloff at frequencies exceeding the inverse DW width, which is also a universal feature.
@@ 2 Theoretical setup
Let us first briefly review basics of DW evolution in the limit $V_{bias}\rightarrow 0$, where the exact $Z_{2}$-symmetry is restored. We assume that the field $\chi$ is initially set to zero cosmologically and starts rolling to its potential minima after the Hubble parameter $H=\dot{a}/a$ drops enough to fulfill the condition $H\lesssim\sqrt{\lambda}v$, from which point on the DW network comes into existence. Note that we neglect any non-gravitational interactions of $\chi$ with the primordial plasma, and hence its thermal mass is set to zero. Soon after its formation, the DW network enters the scaling regime [23]. Numerical simulations [28] suggest that the scaling starts independently of initial conditions at the time $t_{sc}$, when the DW width $\delta_{wall}$, estimated as
@@
becomes small compared to the Hubble radius, i.e., $\delta_{wall}\simeq 0.05H^{-1}_{sc}$. In the scaling regime, the long-range dynamics of the DW network is defined by only one parameter, the expansion rate of the Universe, $H$. In particular, simulations reveal that at any time there is essentially one long DW with a characteristic curvature radius $H^{-1}$ in the Hubble patch. Closed DWs are also present, but their contribution to the overall network is negligible [28]. As a result, the energy density of the network is estimated as
@@
From now on we switch to the case of biased DWs, unless otherwise specified. Conventionally, one defines the DW network annihilation time $\tilde{\tau}_{ann}$ (the reason for the tilde notation will become clear shortly) from the balance between the DW energy density and the potential bias, i.e., $\rho_{wall}\sim V_{bias}$. One also assumes that the scaling of DWs continues up to the annihilation time, so that $\rho_{wall}(\tilde{\tau}_{ann})\sim\sigma_{wall}H_{ann}$. This leads to the dependence $\tilde{\tau}_{ann}\propto 1/\sqrt{\epsilon}$ of so defined annihilation time $\tilde{\tau}_{ann}$ [8]. However, we observe in Sec. 4 that departures from the scaling regime start significantly earlier, thus compromising the physical interpretation of $\tilde{\tau}_{ann}$. Therefore, we consider an alternative way of defining the DW network collapse time, which will prove to be more relevant. Following Refs. [20, 21], we introduce the false vacuum fraction:
@@
Naively one would expect that $\tau_{ann}\sim\tilde{\tau}_{ann}\propto 1/\sqrt{\epsilon}$. However, we will see in Sec. 4 that the simulation data favor a different behavior. Hereafter we adopt the definition of the DW annihilation time (10).
@@ 3 Numerical setup
$\chi^{\prime\prime}+\frac{2}{\tau}\chi^{\prime}-\partial^{2}_{i}\chi+\tau^{2}\chi(\chi^{2}-1)+3\epsilon\tau^{2}\chi^{2}=0\;.$ (15)
@@
Optimally by the end of simulations at the conformal time $\tau_{f}$, the DW width should not be smaller than the lattice spacing, i.e.,
@@
where $N=1024$ or $2048$ is the lattice grid number, $\kappa={\cal O}(1)$ is a constant, $L_{i}=a(\tau_{i})L=L$ is the initial lattice size, so that $L_{i}/N$ is the initial lattice spacing. We take into account that the lattice spacing grows linearly with the scale factor. We also require that the final Hubble horizon is still smaller than the simulation box, i.e.,
@@
Below we fix these final times, but allow for variations of $\kappa^{\prime}$ and $\kappa$. For concreteness, we restrict to three possible choices of $\kappa^{\prime}$ and $\kappa$:
@@
The reason to consider $\kappa^{\prime}/\kappa>1$ is that DW phenomenology and GW signal in particular are dominated by the Hubble scale physics at the latest stages of the network existence. Therefore, we wish the simulation box size to be considerably larger than the Hubble radius at those times, i.e., we allow the ratio $\kappa^{\prime}/\kappa$ to be larger than unity. With such a choice, however, one runs the risk that DWs can become too thin relative to the lattice spacing by the end of simulations. As it has been observed in Ref. [28], this does not invalidate simulations, but leads to the appearance of the artificial features in the UV part of the GW spectrum propagating towards IR with time. The “contamination” is mitigated by choosing the option with $\kappa^{\prime}/\kappa<1$ in Eq. (23).
While we mainly take $\kappa^{\prime}=\pi\kappa/2$ in what follows, in order to reconstruct the GW spectrum, one will need to consider all three options in Eq. (23). In this way, we can “zoom in” the IR, middle, and UV parts of the GW spectrum, and then glue all three parts. The validity of this techniques was proven in Ref. [28], where unbiased DWs have been studied by comparing results obtained with lattices of different resolution.
@@ 4 Numerical results: evolution of domain wall network
Figure 1: Snapshots of DW evolution obtained in the case of dimensionless bias parameter $\epsilon=0.01$. DWs are shown with green, while the regions filled with true and false vacuum are shown with white and red, respectively.
@@
Evolution of the false vacuum fraction demonstrated in Fig. 4 is particularly illuminating. We parameterise the false vacuum fraction as in Refs. [20, 21]:
@@
where $p$ is assumed to be a constant. See also Refs. [31, 32, 33] for theoretical approaches to the problem of DW collapse. Note that the parameter $\tau_{ann}$ entering Eq. (25) matches the definition of the DW annihilation time given in Eq. (10). Compared to Ref. [21], we observe that the power $p$ may vary as a function of the potential bias, and at this point we observe dependence on the initial conditions. Namely, the variation is rather strong for $k_{cut}=1$ and $k_{cut}=5$, while for $k_{cut}=0.3$ the parameter $p$ gets confined to a narrow range near $p=3.3$. In this “stabilised” regime, values of the parameter $p$ are in a good agreement with those of Ref. [21].
@@
where $C(k_{cut})$ is a coefficient, which is independent of $\epsilon$, but slightly varying with the momentum cutoff $k_{cut}$. In particular, one has $C(1)\approx 1$, $C(0.3)\approx 1.4$, and $C(5)\approx 0.9$. Recalling that the initial time $\tau_{i}$ is chosen in such a way that $H_{i}=1/(a_{i}\tau_{i})=\sqrt{\lambda}v$, one can rephrase Eq. (26) in terms of the Hubble rate at radiation domination as
@@
$\frac{H_{ann}}{\sqrt{\lambda}v}\approx\frac{0.1}{C^{2}(k_{cut})}\cdot\left(\frac{\epsilon}{\lambda v}\right)^{2/3}\;.$ (27)
@@
Interestingly, the dependence $\tau_{ann}(\epsilon)$ as in Eq. (26) is consistent with the standard estimate $\rho_{wall}\sim V_{bias}$ provided that the DW energy density drops as $\rho_{wall}\propto 1/a^{3}$ rather than $\rho_{wall}\propto 1/a^{2}$. In other words, in order to satisfy the standard estimate $\rho_{wall}\sim V_{bias}$, the DW density must decrease faster than in the scenario with unbiased DWs. This correlates with the behavior of the scaling parameter shown in Fig. 3 suggesting that the potential bias starts impacting DWs early in their evolution. Such a similarity of the behavior of biased DWs with the case of non-interacting dust particles might be not a coincidence (e.g., it may be partly due to growing amount of closed DW behaving like dust inside a Hubble patch), and it is worth investigating in the future. Notably, the observed behavior $\tau_{ann}(\epsilon)$ is largely independent of initial conditions. Namely, unlike the parameter $p$ and the coefficient $C(k_{cut})$, the relation $\tau_{ann}\propto 1/\epsilon^{1/3}$ is robust against variations of the momentum cutoff $k_{cut}$. In other words, we find that DW annihilation time is parametrically smaller (with respect to $\epsilon$) than naively expected. As we will see in the next section, this leads to drastic decrease of GW power as compared to the commonly assumed estimates.
@@
To derive the relation (26) we have probed the ranges $\epsilon=[0.0035,0.05]$ and $\epsilon=[0.01,0.05]$ using $2048^{3}$ and $1024^{3}$ lattices, respectively. Note that the similar conclusion can be derived from the analysis of false vacuum fraction performed in Ref. [21], where the narrow range $\epsilon=[4\cdot 10^{-4},8\cdot 10^{-4}]$ has been explored with higher resolution simulations. One can check that the behavior (26) is indeed obeyed except for the smallest $\epsilon$ there. Steepening of the behavior $\tau_{ann}(\epsilon)$ in the range of very small $\epsilon$ is likely attributed to the fact that in this case the network decay takes place at sufficiently late times, when effects, related to finite lattice size/resolution, become significant.
@@
Figure 6: Comparison of the potential bias impact on DW evolution caused by the symmetry breaking potentials $V_{breaking}=\epsilon\chi^{3}$, $\epsilon\chi v^{2}$, and $\epsilon(\chi^{5}/v^{2}+\chi^{6}/v^{3})$. The bias parameter $\epsilon$ has been set to $\epsilon=0.05$ (in units of $\lambda v$). The impact is demonstrated in terms of the scaling parameter $\xi$ (left panel) and the false vacuum fraction ${\cal F}_{fv}$ (right panel). Simulations have been carried out with $1024^{3}$ lattice.
@@
Let us consider other choices of $Z_{2}$-symmetry breaking caused by the linear term $V_{breaking}=\epsilon\chi v^{2}$ as well as the non-renormalisable term of the form $V_{breaking}=\epsilon\chi^{5}/v^{2}+\epsilon^{\prime}\chi^{6}/v^{3}$. In the latter case, the term $\epsilon^{\prime}\chi^{6}/v^{3}$ with some positive constant $\epsilon^{\prime}$ has been added to keep the overall potential bounded from below. As it is clear from Fig. 6, the impact of the higher order potential bias on DW evolution is almost indistinguishable from that of the cubic bias. On the other hand, the effect of the linear bias is dramatically different. The likely reason is that the linear bias term changes behavior of its potential $V(\chi)$ around the minimum $\chi=0$ leading to non-zero $\partial V/\partial\chi$ at $\chi=0$. As a result, one of the minima becomes slightly preferred meaning that the linear potential bias triggers the population bias. This may explain, why the DW network is destroyed at much faster rate for the same $\epsilon$ [34], and consequently the expected GW signal is suppressed relative to the case of the cubic bias. In this work we focus on the effects, which are entirely due to the potential bias, and omit any further discussion of the case $V_{breaking}\propto\chi$.